OpenStudy (anonymous):

What is 5^-3 kg expressed as a quotient?

OpenStudy (ybarrap):

The negative exponent means put inverse of \(5^3\). So now what do you think this looks like?

OpenStudy (ybarrap):

$$ \large{ 5^{-3}=\frac{1}{5^3} } $$ That's what a negative exponent means.

OpenStudy (triciaal):

other approach/ more details the quotient is what you get when you divide 5^-3 means 5^0-3 when you have the same base and subtract exponents you are dividing 5^0-3 = 5^0 / 5^3 and 5^0 = 1 thus 5^-3 = 1/5^3 as above